OpenStudy (anonymous):

What is the solution to the equation? 6x – 27 = –3 A. x = –5 B. x = –4 C. x = 4 D. x = 5

OpenStudy (anonymous):

Fan & Medal will be given

OpenStudy (anonymous):

6x -27 = -3 you need to get x on its own, so 6x -27 = -3 +27 +27 _______________ 6x = 24 divide by 6 to get x alone x = 4, so C hope this helps!